Circular No. 9105

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2010 A2 (LINEAR) An apparently asteroidal object reported by the LINEAR project (dis-

covery observation tabulated below) and placed on the NEO Confirmation Page has been reported as cometary and difficult to measure by several CCD astrometrists. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector, Jan. 7.0) found the object to be a headless comet with a straight tail in individual 20-s exposures (in a 2600-s integration, the tail has a length of at least 130′′ in p.a. 280^{o} and a width of 10′′–14′′ along its entire length — brightest in a 20′′ area starting 6′′ from the eastern end, fading gradually to the west); there is no obvious central condensation, but the center of brightness may be ≈ 11′′ from the eastern end, and the positional measurement was very difficult. On images obtained during Jan. 7.84–7.87, Birtwhistle found a similar appearance to the previous night but found the tail to be at least 260′′ long in p.a. 279^{o}. W. H. Ryan (Magdalena Ridge Observatory, 2.4-m f/8.9 reflector, Jan. 7.20–7.29) reports an ill-defined elongated nuclear condensation and a very distinct, bright tail in p.a. ∼ 280^{o}. H. Sato (Ota-ku, Tokyo, Japan, remotely using a 0.25-m f/3.4 reflector of the RAS Observatory, near Mayhill, NM, U.S.A., Jan. 7.2) notes a very diffuse coma 10′′ in diameter and a 2′.2 tail toward p.a. 280^{o}. On Jan. 7.3, E. C. Beshore (Mt. Lemmon Survey, 1.5-m reflector) followed the tail out to nearly 4′ in p.a. 285^{o}. K. Kadota (Ageo, Saitama-ken, Japan, 0.25-m f/5 reflector, Jan. 7.7) found the comet to be diffuse, with a weak central condensation and a faint coma of diameter 0′.5, showing a straight, narrow tail 5′ long in p.a. 280^{o}.

2010 UT α_{2000} δ_{2000} Mag. Observer
Jan. 6.27428 7^{h}10^{m}16.^{s}81 +28^{o}21′35′′.1 19.9 LINEAR

Further astrometry, the following very preliminary elliptical orbital elements, and an ephemeris are given on MPEC 2010-A32.

T = 2010 May 7.565 TT  ω = 227.^{o}300
e = 0.63483  Ω = 313.525 }2000.0
q = 1.35759 AU  i = 6.068
a = 3.71766 AU n^{o} = 0.137499 P = 7.17 years

COMET C/2007 Q3 (SIDING SPRING) Visual total-magnitude estimates by M. Meyer, Limburg, Germany

(15-cm refractor): Dec. 21.19 UT, 9.7; 26.21, 9.7; 29.20, 9.6.

2010 January 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green